OpenStudy (anonymous):

solve the system by addition,if the system is inconsistent and has no solution,state this .If the system is dependent,write the form pf the solution for any real number x. 8x-3y=34 -2+9y=8

OpenStudy (anonymous):

I am solving it first so dont give answer yet until I solve mine thanks

OpenStudy (anonymous):

I got (48.7,4.9)

OpenStudy (anonymous):

No solution it is inconsistent check it for me if am correct

OpenStudy (anonymous):

that is -2x or -2 only -2+9y=8 9y=10 y=10/9 put in first get x

OpenStudy (anonymous):

consistent 8x-3y=34 (1) -2x+9y=8 (2) do this (1)+fourtimes(2)

OpenStudy (anonymous):

x=14/3 and y=10/9